# GP-GS: Gaussian Processes for Enhanced Gaussian Splatting

### v3

@@ Abstract
3D Gaussian Splatting has emerged as an efficient photorealistic novel view synthesis method. However, its reliance on sparse Structure-from-Motion (SfM) point clouds consistently compromises the scene reconstruction quality. To address these limitations, this paper proposes a novel 3D reconstruction framework Gaussian Processes Gaussian Splatting (GP-GS), where a multi-output Gaussian Process model is developed to achieve adaptive and uncertainty-guided densification of sparse SfM point clouds. Specifically, we propose a dynamic sampling and filtering pipeline that adaptively expands the SfM point clouds by leveraging GP-based predictions to infer new candidate points from the input 2D pixels and depth maps. The pipeline utilizes uncertainty estimates to guide the pruning of high-variance predictions, ensuring geometric consistency and enabling the generation of dense point clouds. The densified point clouds provide high-quality initial 3D Gaussians to enhance reconstruction performance. Extensive experiments conducted on synthetic and real-world datasets across various scales validate the effectiveness and practicality of the proposed framework. Code is available.
@@
Figure 1: Illustration of the point cloud densification and improved rendering results. (a) Conventional SfM only yields sparse point clouds, which poorly capture rich scene details. Our Multi-Output Gaussian Process model addresses this issue by densifying the sparse point clouds across various datasets, achieving 71% (indicated by $R^{2}$ score (Edwards et al. 2008)) position and colour accuracy within the predicted confidence interval. (b) Rendered images comparison with 3DGS* (official improvements). GP-GS substantially enhances rendering quality, particularly in intricate regions like dense foliage or under challenging lighting conditions.
1 Introduction
@@
Novel View Synthesis (NVS) is a fundamental yet challenging problem in computer vision and computer graphics, aiming to generate novel viewpoints of a given scene from multi-view observations. It plays a critical role in applications such as Digital Twinning (Wang et al. 2024b), virtual reality (Fei et al. 2024), and robotics (Xiong et al. 2024; Wang et al. 2024a). Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) have revolutionized NVS by implicitly modelling volumetric scene representations, achieving high-fidelity rendering without explicit 3D geometry reconstruction. However, NeRF-based methods suffer from computational inefficiency and the requirement of dense sampling along rays, leading to slow inference speeds despite recent acceleration efforts (Guo and Wang 2024; Mildenhall et al. 2021; Roessle et al. 2022; Zhang et al. 2020),.
@@
To overcome these limitations, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has emerged as a promising alternative for real-time rendering. Unlike NeRF, which relies on ray-marching and volumetric integration, 3DGS represents scenes explicitly using a set of 3D Gaussians with learnable attributes like positions and colours. The rasterization-based splatting strategy can be considered a forward rendering technique that avoids costly ray sampling and enables efficient parallel rendering, making it a compelling choice for NVS applications. However, existing 3DGS pipelines rely heavily on Structure-from-Motion (SfM) to extract 3D points by detecting and matching texture features like SIFT points, leading to sparse or incomplete point clouds, which is particularly the case in texture-less or clustered regions, leading to poor initialization of 3D Gaussians. Consequently, artefacts arise in the rendering process, causing a loss of fine details. To mitigate the limitations of sparse SfM-based reconstructions, the Adaptive Density Control (ADC) strategy has been introduced to 3DGS pipelines. ADC aims to improve scene coverage by dynamically duplicating Gaussians in underrepresented regions and removing redundant ones in over-reconstructed areas. However, ADC operates without explicit geometric priors, often producing noisy distributions that fail to adhere to the underlying scene structure. As a result, this leads to blurry reconstructions and poor occlusion handling, limiting the overall fidelity of 3DGS-based novel view synthesis. These challenges significantly limit the fidelity and robustness of 3DGS-based NVS methods, motivating the community for a more structured and adaptive approach to point cloud refinement and rendering.
@@
In this paper, we propose GP-GS (Gaussian Processes for Enhanced 3D Gaussian Splatting), a novel framework designed to enhance the initialization of 3D Gaussians and improve rendered quality, especially in complex regions with densely packed objects (e.g., foliage) or under challenging lighting conditions. To be specific, we introduce a Multi-Output Gaussian Process (MOGP) model to achieve adaptive and uncertainty-guided densification of sparse SfM point clouds. Specifically, we formulate point cloud densification as a continuous regression problem, where the MOGP learns a mapping from 2D image pixels and their depth priors to a denser point cloud with position and colour information. To ensure a structured and robust densification process, we propose an adaptive neighbourhood-based sampling strategy, where pixels are selected as candidate MOGP inputs for densification. For each sampled pixel, we predict its corresponding 3D point cloud attributes (position, colour, variance) using the MOGP model. To further refine the densified point cloud, we apply a variance-based filtering method, which removes high-uncertainty predictions, thereby reducing noise accumulation and preserving high-confidence reconstructions. The densified point clouds provide high-quality 3D Gaussian blobs (or Gaussians for brevity) to enhance the reconstruction performance.
@@
Our contributions can be summarized as: 1) Gaussian Processes for Point Cloud Densification: We propose a MOGP model to densify sparse SfM point clouds by learning mappings from 2D image pixels and depth information to 3D positions and colours, with uncertainty awareness. 2) Adaptive Sampling and Uncertainty-Guided Filtering: We introduce an adaptive neighbourhood-based sampling strategy that generates candidate inputs for MOGP for 3D points prediction, followed by variance-based filtering to remove high uncertainty predictions, ensuring geometric consistency and enhancing rendering quality. 3) Our GP-GS framework can be seamlessly integrated into existing SfM-based pipelines, making it a flexible plug-and-play module that improves the rendering quality of other NVS models.
2 Related Work
2.1 Gaussian Processes
@@
Gaussian Processes (GPs) are a collection of random variables, any finite number of which subjects to a joint Gaussian distribution (Rasmussen 2003). GPs are particularly effective in handling sparse data, making them well-suited for scenarios with limited observations (Wang et al. 2021). Their flexibility and built-in uncertainty quantification enable robust predictions, which has led to their widespread adoption in various computer vision tasks (Lu et al. 2023; Zhou et al. 2023; Zhu, Wang, and Mihaylova 2021; Yasarla, Sindagi, and Patel 2020).
@@
Mathematically, a GP is fully specified by its mean function $m(\mathbf{x})$ and covariance function $k\left(\mathbf{x},\mathbf{x}^{\prime}\right)$, a.k.a. kernel function, where $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are inputs. The covariance function typically depends on a set of hyperparameters $\varphi$, such as length scale $l$, variance $\sigma^{2}$, or other parameters depending on the kernel type (e.g., Matérn). A GP is usually used as a prior over a latent function defined as:
@@
Training a GP model involves optimizing the hyperparameters $\theta$ of the kernel function to maximize the likelihood of the observed data. This is achieved by maximizing the log marginal likelihood, given by:
$\log p(\mathbf{y}\mid\mathbf{X},\theta)=-\frac{1}{2}\mathbf{y}^{T}\mathbf{K}^{-1}\mathbf{y}-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log 2\pi,$ (2)
where $\mathbf{y}$ are the observed outputs, $\mathbf{X}$ are the observed inputs, $\mathbf{K}\in\mathbb{R}^{n\times n}$ is the Gram matrix with entries computed using the kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j})$ across all training inputs, and $n$ is the number of training points. After training, the predictive distribution at a new test point $\mathbf{x}_{*}$ follows a Gaussian distribution: $f\left(\mathbf{x}_{*}\right)\mid\mathbf{X},\mathbf{y},\mathbf{x}_{*}\sim\mathcal{N}\left(\mu_{*},\sigma_{*}^{2}\right)$, where the predictive mean and variance are computed as:
@@
where $\mathbf{k}_{*}=[k(\mathbf{x}_{*},\mathbf{x}_{i})]\in\mathbb{R}^{n\times 1}$ represents the covariance between the test point $\mathbf{x}_{*}$ and all training points, while $k(\mathbf{x}_{*},\mathbf{x}_{*})$ denotes the self-variance of the test point.
2.2 3D Gaussian Splatting
@@
3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) employs a forward rendering approach. This technique represents scenes using 3D Gaussians and achieves efficient rendering by directly projecting these Gaussians onto the 2D image plane. By circumventing the need for extensive ray sampling and complex volumetric integration, 3DGS facilitates real-time, high-fidelity scene reconstruction. Specifically, it computes pixel colours by depth sorting and $\alpha$-blending of projected 2D Gaussians. This method avoids the complex calculation of ray marching and volume integration and can achieve real-time high-quality rendering and NVS. Several works have enhanced 3DGS, such as mitigating artefacts arising from camera pose sensitivity (Yu et al. 2024). Others manage points to improve rendering quality (Yang et al. 2024a; Zhang et al. 2024; Bulò, Porzi, and Kontschieder 2024). To the best of our knowledge, only a few studies have explored the densification of SfM point clouds for 3DGS performance improvement (Cheng et al. 2024; Chan et al. 2024). While there is evidence showing that densification helps to improve the rendering quality of 3DGS, they often overlook the mathematical relationship between pixels and point clouds derived from the initial sparse SfM.
To bridge this gap, our method introduces a MOGP model, to learn the mapping relationship between pixels and point clouds, which facilitates the densification of 3D Gaussians, thereby enhancing the rendering quality of 3DGS.
@@
Figure 2: The overview of GP-GS. (a) Multi Views and Depth Priors: Multi-view images are first processed with depth estimating models like Depth Anything (Yang et al. 2024b) to extract per-view depth maps. (b) Point Cloud Densification: Sparse point cloud is initially reconstructed using SfM. Next, MOGP is trained to take pixel coordinates and depth as inputs $\mathbf{X}=(u,v,d)$ and predicts dense point clouds $\mathbf{Y}=(x,y,z,r,g,b)$ with uncertainty (variance). The loss function ensures an optimal mapping between input pixels and point clouds. (c) Uncertainty-Based Filtering: The predicted dense point clouds undergo filtering based on uncertainty. This improves the quality of the dense point clouds that are used to initialize dense 3D Gaussians, which are then optimized to refine geometric details. The final rendered novel views demonstrate the effectiveness of GP-GS in reconstructing fine details while maintaining structural coherence.
3 Preliminaries
@@
3DGS (Kerbl et al. 2023) initialize SfM sparse point clouds as 3D Gaussians (ellipsoid shaped), each defined by specific parameters such as position (mean), rotation, scales, covariance, opacity $\alpha$, and colour represented as spherical harmonics. For the sake of brevity, this paper abuses $\mathbf{x}$, representing the position of Gaussians, to define 3D Gaussian $G$ as follows:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})},$ (4)
@@
where $\bm{\mu}\in\mathbb{R}^{3\times 1}$ is the mean vector, $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ is the covariance matrix. The 3D covariance $\mathbf{\Sigma}$ is a positive semi-definite matrix, which can be denoted as: $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}\in\mathbb{R}^{3\times 3}$ is an orthogonal rotation matrix, and $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a diagonal scale matrix.
@@
3D Gaussians will be projected to 2D image space using the splatting-based rasterization technique (Zwicker et al. 2001b). Specifically, the transformation is approximated with a first-order Taylor expansion at a projected point in the camera coordinate frame. This ensures efficient and accurate rendering from a given viewpoint. The 2D covariance matrix $\bm{\Sigma}^{\prime}$, which describes the elliptical shape of each Gaussian in the image space, is then computed as:
$\bm{\Sigma}^{\prime}=\mathbf{JW}\bm{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T},$ (5)
@@
where $\mathbf{J}$ is the Jacobian of the affine approximation of the projective transformation, and $\mathbf{W}$ denotes the view transformation matrix (Zwicker et al. 2001a). The colour of each pixel is calculated by blending sorted Gaussians based on $\alpha$, as follows:
@@
where $n$ is the number of points, $c_{i}$ is the color of the $i-th$ point, $\alpha_{i}$ can be obtained by evaluating a projected 2D Gaussian with covariance $\bm{\Sigma}^{\prime}$ multiplied with a learned opacity for each point.
4 Methodology
4.1 Overview
This paper proposes a novel framework GP-GS that enhances 3D Gaussians initialization, thereby improving 3DGS rendering quality. The overview of GP-GS is presented in Figure 2. First, addressing the limitations of sparse SfM point clouds, we propose the MOGP model to formulate point cloud densification as a regression problem. (Sec. 4.2). The training strategy of MOGP is introduced in Appendix A.1. Next, we adaptively densify sparse SfM point clouds and propose an uncertainty-based filtering strategy to remove high-uncertainty predictions. (Sec. 4.3).
4.2 Multi Output Gaussian Process

### v4

@@ Abstract
3D Gaussian Splatting has emerged as an efficient photorealistic novel view synthesis method. However, its reliance on sparse Structure-from-Motion (SfM) point clouds often limits scene reconstruction quality. To address the limitation, this paper proposes a novel 3D reconstruction framework, Gaussian Processes enhanced Gaussian Splatting (GP-GS), in which a multi-output Gaussian Process model is developed to enable adaptive and uncertainty-guided densification of sparse SfM point clouds. Specifically, we propose a dynamic sampling and filtering pipeline that adaptively expands the SfM point clouds by leveraging GP-based predictions to infer new candidate points from the input 2D pixels and depth maps. The pipeline utilizes uncertainty estimates to guide the pruning of high-variance predictions, ensuring geometric consistency and enabling the generation of dense point clouds. These densified point clouds provide high-quality initial 3D Gaussians, enhancing reconstruction performance. Extensive experiments conducted on synthetic and real-world datasets across various scales validate the effectiveness and practicality of the proposed framework. The code is available at https://github.com/zhihaohaoran/GPGS.
@@ I Introduction
Figure 1: Illustration of our MOGP-based enhancement for 3D Gaussian Splatting. (a) Conventional SfM often yields sparse point clouds, which struggle to capture fine scene details. Our Multi-Output Gaussian Process model mitigates this by densifying the sparse point clouds across various datasets, achieving 71% accuracy in position and colour prediction (as indicated by the $R^{2}$ score [7]) within the predicted confidence interval. (b) Compared to 3DGS* (official improvements), our approach (GP-GS) generates higher-fidelity renderings, particularly in challenging regions such as dense foliage and shaded areas. (c) Quantitative results across multiple datasets show that GP-GS achieves a 0.09 dB gain over 3DGS* and a 0.17 dB gain over Scaffold-GS on Mip-NeRF 360, and consistently outperforms baselines across all datasets.
@@
Novel View Synthesis (NVS) is a fundamental yet challenging problem in computer vision and computer graphics, aiming to generate novel viewpoints of a given scene from multi-view observations. It plays a critical role in applications such as digital twinning [29], virtual reality [9, 33], and robotics [27, 32]. Neural Radiance Fields (NeRF) [19] have revolutionized NVS by implicitly modelling volumetric scene representations, achieving high-fidelity rendering without explicit 3D geometry reconstruction. However, NeRF-based methods suffer from computational inefficiency and the requirement of dense sampling along rays, leading to slow inference speeds despite recent acceleration efforts [19, 23, 40].
@@
To overcome these limitations, 3D Gaussian Splatting (3DGS) [13] and its variants [17, 12] have emerged as promising alternatives for real-time rendering. Unlike NeRF, which rely on ray-marching and volumetric integration, 3DGS represents scenes explicitly using a set of 3D Gaussians with learnable attributes like positions and colours. The rasterization-based splatting strategy can be considered a forward rendering technique that avoids costly ray sampling and enables efficient parallel rendering, making it a compelling choice for NVS applications. However, existing 3DGS pipelines rely heavily on Structure-from-Motion (SfM) to extract 3D points by detecting and matching texture features like SIFT (Scale-Invariant Feature Transform) points, leading to sparse or incomplete point clouds, which is particularly the case in texture-less or clustered regions, respectively, leading to poor initialization of 3D Gaussians. Consequently, artefacts arise in the rendering process, causing a loss of fine details. To mitigate the limitations of sparse SfM-based reconstructions, the Adaptive Density Control (ADC) strategy has been introduced to 3DGS pipelines. ADC aims to improve scene coverage by dynamically duplicating Gaussians in underrepresented regions and removing redundant ones in over-reconstructed areas. However, ADC operates without explicit geometric priors, often producing noisy distributions that fail to adhere to the underlying scene structure. As a result, this leads to blurry reconstructions and poor occlusion handling, limiting the overall fidelity of 3DGS-based NVS. These challenges significantly limit the fidelity and robustness of 3DGS-based NVS methods, motivating the community for a more structured and adaptive approach to point cloud refinement and rendering.
@@
In this paper, we propose GP-GS (Gaussian Processes for Enhanced 3D Gaussian Splatting), a novel framework designed to enhance the initialization of 3D Gaussians and improve rendered quality, especially in complex regions with densely packed objects (e.g., foliage) or under challenging lighting conditions. At the core of GP-GS is a Multi-Output Gaussian Process (MOGP) model that enables adaptive, uncertainty-guided densification of sparse point clouds generated by Structure-from-Motion (SfM). We formulate this task as a continuous regression problem, where the MOGP learns a mapping from 2D image pixels and their associated depth priors to a denser point cloud enriched with 3D position and color information. To ensure a structured and robust densification process, we propose an adaptive neighbourhood-based sampling strategy, where pixels are selected as candidate MOGP inputs for densification. For each sampled pixel, we predict its corresponding 3D point cloud attributes (position, colour, variance) using the MOGP model. To further refine the densified point cloud, we apply variance-based filtering to remove high-uncertainty predictions, reducing noise accumulation and preserving high-confidence reconstructions. The densified point clouds ultimately provide high-quality 3D Gaussian blobs (or Gaussians for brevity) to enhance the reconstruction performance.
@@
Our contributions can be summarized as: 1) Gaussian Processes for Point Cloud Densification: We propose a MOGP model to densify sparse SfM point clouds by learning mappings from 2D image pixels and depth information to 3D positions and colours, with uncertainty awareness. 2) Adaptive Sampling and Uncertainty-Guided Filtering: We introduce an adaptive neighbourhood-based sampling strategy that generates candidate inputs for MOGP for 3D points prediction, followed by variance-based filtering to remove high uncertainty predictions, ensuring geometric consistency and enhancing rendering quality. 3) Our GP-GS framework can be seamlessly integrated into existing SfM-based pipelines, serving as a flexible plug-and-play module that improves the rendering quality of other NVS models.
II Related Work
II-A Gaussian Processes
@@
Gaussian Processes (GPs) are a collection of random variables, any finite number of which are subject to a joint Gaussian distribution [21]. GPs are particularly effective in handling sparse data, making them well-suited for scenarios with limited observations [28, 25]. Their flexibility and built-in uncertainty quantification enable robust predictions and uncertainty-informed signal processing, which has led to their widespread adoption in various computer vision tasks [43, 24].
@@
Mathematically, a GP is fully specified by its mean function $m(\mathbf{x})$ and covariance function $k\left(\mathbf{x},\mathbf{x}^{\prime}\right)$, a.k.a. kernel function, where $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are inputs. The kernel function typically depends on a set of hyperparameters $\boldsymbol{\theta}$, such as length scale $l$, signal variance $\sigma_{f}^{2}$, or other parameters depending on the kernel type (e.g., Matérn). A GP is usually used as a prior over a latent function defined as:
@@
Training a GP model involves optimizing the hyperparameters $\boldsymbol{\theta}$ of the kernel function to maximize the likelihood of the observed data. This is achieved by maximizing the log marginal likelihood, given by:
$\log p(\mathbf{y}\mid\mathbf{X},\boldsymbol{\theta})=-\frac{1}{2}\mathbf{y}^{T}\mathbf{K}^{-1}\mathbf{y}-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log 2\pi,$ (2)
where $\mathbf{y}$ are the observed outputs, $\mathbf{X}$ are the observed inputs, $\mathbf{K}\in\mathbb{R}^{n\times n}$ is the Gram matrix with entries computed using the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ across all training inputs, and $n$ is the number of training points. After training, the predictive distribution at a new test point $\mathbf{x}_{*}$ follows a Gaussian distribution: $f\left(\mathbf{x}_{*}\right)\mid\mathbf{X},\mathbf{y},\mathbf{x}_{*}\sim\mathcal{N}\left(\mu_{*},\sigma_{*}^{2}\right)$, where the predictive mean and variance are computed as:
@@
where $\mathbf{k}_{*}=[k(\mathbf{x}_{*},\mathbf{x}_{i})]^{T},i=1,\cdots,n,\in\mathbb{R}^{n\times 1}$ represents the covariance between the test point $\mathbf{x}_{*}$ and all training points, while $k(\mathbf{x}_{*},\mathbf{x}_{*})$ denotes the variance of the test point itself.
II-B Neural Radiance Fields
NeRF [19] utilizes implicit representation for the first time to achieve photo-realistic perspective synthesis, subsequently inferring the 3D structure of the scene. It utilizes a limited number of input views to train a neural network to represent a continuous volumetric scene, enabling the generation of new perspectives of the scene once the neural network is trained. To be specific, given the 3D location $\mathbf{p}=[x,y,z]^{T}$ of a spatial sampling point (to be rendered) and 2D view direction $\mathbf{d}=[\theta,\phi]^{T}$ of the camera, NeRF predicts the color $\mathbf{c}=[r,g,b]^{T}$ of the sampling point and the volume density $\sigma$ through a multi-layer perceptron (MLP) neural network [20] which can be represented as $F_{\Theta}:(\mathbf{p},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$, where $F_{\Theta}$ is the MLP parameterized by $\Theta$. The color is estimated by volumetric rendering via quadrature, which can be formulated as
[…]
II-C 3D Gaussian Splatting
@@ II-B Neural Radiance Fields
In contrast to NeRF’s inverse rendering methodology, 3D Gaussian Splatting (3DGS) [13] employs a forward rendering approach. This technique represents scenes using 3D Gaussians and achieves efficient rendering by directly projecting these Gaussians onto the 2D image plane. By circumventing the need for extensive ray sampling and complex volumetric integration, 3DGS facilitates real-time, high-fidelity scene reconstruction. Specifically, it computes pixel colours by depth sorting and $\alpha$-blending of the projected 2D Gaussians. This method avoids the complex calculation of ray marching and volume integration and can achieve real-time high-quality rendering and NVS. Several works have enhanced 3DGS, such as mitigating artifacts arising from camera pose sensitivity [38]. Others manage points to improve rendering quality [35, 42, 2, 18]. To the best of our knowledge, only a few studies have explored the densification of SfM point clouds for 3DGS performance improvement [6, 3]. To address this gap, we introduce a probabilistic model based on MOGP, which learns a continuous mapping from image pixels and depth priors to dense 3D point clouds with associated colour attributes. This approach not only improves reconstruction fidelity from sparse or dense SfM inputs but also provides principled uncertainty estimates, enabling adaptive pixel sampling and variance-based filtering for enhanced rendering and NVS.
While recent SfM-free methods such as HGM [4] and InstantSplat [8] achieve impressive results using pre-trained priors (e.g., MASt3R [15]) to initialize geometry and camera poses, they are highly dependent on the quality of learned priors and often require co-visibility filtering to reduce redundancy. In contrast, our method builds upon the widely used SfM pipeline, which offers a flexible, interpretable, and data-driven foundation without relying on large-scale pretrained models. Therefore, we focus our comparative analysis on SfM-based methods, aiming to improve this practical and widely adopted pipeline. We view SfM-free approaches as complementary but operating under different assumptions and problem settings.
@@
Figure 2: The overview of GP-GS. (a) Key Frame and Depth Priors: Sparse point cloud is initially obtained using SfM. We then perform pixel-to-point cloud matching and select the key frame that contributes best to the point cloud. The selected key frame is processed with depth estimation models like Depth Anything [36] to extract the depth map. (b) Point Cloud Densification: MOGP is trained to take key frame pixel coordinates and depth as inputs $\mathbf{X}=[u,v,d]$ and predicts dense point clouds $\mathbf{Y}=[x,y,z,r,g,b]$ with uncertainty (variance). The loss function ensures an optimal mapping between input pixels and point clouds. (c) MOGP Testing: The adaptive sampling strategy adaptively generate sampling pixels within the image domain, and uncertainty-based filtering removes unreliable predictions. This results in a refined dense point cloud. (d) MOGP Output: The sparse and the predicted point clouds are merged and then used to initialize 3D Gaussians, which are then optimized to refine geometric details. The final rendered novel views demonstrate the effectiveness of GP-GS in reconstructing fine details while maintaining structural coherence.
III Preliminaries
@@
3DGS [13] initialize SfM sparse point clouds as 3D Gaussians (ellipsoid shaped), each defined by specific parameters such as position (mean), rotation, scales, covariance, opacity $\alpha$, and colour represented as spherical harmonics. For the sake of brevity, this paper abuses $\mathbf{x}$, representing the position of Gaussians, to define 3D Gaussian $G$ as follows:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})},$ (5)
@@
where $\boldsymbol{\mu}\in\mathbb{R}^{3\times 1}$ is the mean vector, $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$ is the covariance matrix. The 3D covariance $\mathbf{\Sigma}$ is a positive semi-definite matrix, which can be denoted as: $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}\in\mathbb{R}^{3\times 3}$ is an orthogonal rotation matrix, and $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a diagonal scale matrix.
@@
3D Gaussians will be projected to 2D image space using the splatting-based rasterization technique. Specifically, the transformation is approximated with a first-order Taylor expansion at a projected point in the camera coordinate frame. This ensures efficient and accurate rendering from a given viewpoint. The 2D covariance matrix $\boldsymbol{\Sigma}^{\prime}$, which describes the elliptical shape of each Gaussian in the image space, is then computed as:
$\boldsymbol{\Sigma}^{\prime}=\mathbf{JW}\boldsymbol{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T},$ (6)
@@
where $\mathbf{J}$ is the Jacobian of the affine approximation of the projective transformation, and $\mathbf{W}$ denotes the view transformation matrix [44]. The colour of each pixel is calculated by blending sorted Gaussians based on $\alpha$, as follows:
@@
where $n$ is the number of points, $c_{i}$ is the color of the $i-th$ point, $\alpha_{i}$ can be obtained by evaluating a projected 2D Gaussian with covariance $\boldsymbol{\Sigma}^{\prime}$ multiplied with a learned opacity for each point.
IV Methodology
IV-A Overview
This paper proposes a novel framework GP-GS that enhances 3D Gaussians initialization, thereby improving 3DGS rendering quality. The overview of GP-GS is presented in Figure 2. First, addressing the limitations of sparse SfM point clouds, we propose the MOGP model to formulate point cloud densification as a regression problem. (Sec. IV-B). The training strategy of MOGP is introduced in Algorithm 1. Next, we adaptively densify sparse SfM point clouds and propose an uncertainty-based filtering strategy to remove high-uncertainty predictions. (Sec. IV-C).
IV-B Multi Output Gaussian Process